OpenStudy (anonymous):

y' - y = 2te^(2t) ; y(0) = 1 first order differential equation.. i need to know how to solve it well basic steps i know it is following the general form dy/dt = p(t)y = g(t)

OpenStudy (anonymous):

got it nevermind :)

OpenStudy (jamesj):

There is a 'particular' solution, yp, and the homogeneous solution yh. The solution yh satisfies the equation y' - y = 0, and the general solution y = yp + C.yh satisfies the inhomogeneous equation of the problem So find these two solutions, add them up, and then make them meet the initial condition. Step 1: Homogeneous solution if y' - y = 0, you can separate variables and dy/y = dx. Thus yh = .... Step 2: Particular solution There is a definite procedure for this that I know from more general theory, which I won't give you here and I'm guessing you haven't learnt it yet. But if you had to guess, you would say it's something like, yp(x) = Ate^(2t) + Be^(2t). Put this into the equation and you'll find what A and B work. Step 3: Solve the IVP Write now the general solution y(x) = yp(x) + C.yh(x) and substitute the initial condition y(0) = 1 to solve for C.

OpenStudy (anonymous):

thanks jamesj i think i get it now i was making a slight error in my initial value so my factor wasn't right